# Quantum-Brain: Quantum-Inspired Neural Network Approach to Vision-Brain Understanding

### v1

@@ Abstract
Vision-brain understanding aims to extract semantic information about brain signals from human perceptions. Existing deep learning methods for vision-brain understanding are usually introduced in a traditional learning paradigm missing the ability to learn the connectivities between brain regions. Meanwhile, the quantum computing theory offers a new paradigm for designing deep learning models. Motivated by the connectivities in the brain signals and the entanglement properties in quantum computing, we propose a novel Quantum-Brain approach, a quantum-inspired neural network, to tackle the vision-brain understanding problem. To compute the connectivity between areas in brain signals, we introduce a new Quantum-Inspired Voxel-Controlling module to learn the impact of a brain voxel on others represented in the Hilbert space. To effectively learn connectivity, a novel Phase-Shifting module is presented to calibrate the value of the brain signals. Finally, we introduce a new Measurement-like Projection module to present the connectivity information from the Hilbert space into the feature space. The proposed approach can learn to find the connectivities between fMRI voxels and enhance the semantic information obtained from human perceptions. Our experimental results on the Natural Scene Dataset benchmarks illustrate the effectiveness of the proposed method with Top-1 accuracies of $95.1\%$ and $95.6\%$ on image and brain retrieval tasks and an Inception score of $95.3\%$ on fMRI-to-image reconstruction task. Our proposed quantum-inspired network brings a potential paradigm to solving the vision-brain problems via the quantum computing theory.
Figure 1: Quantum-Brain for Vision-Brain Understanding. Motivated by the connectivities in brain signals and entanglement in quantum theory, we propose Quantum-Brain, a quantum-inspired neural network to explore the connectivities in brain signals for vision-brain understanding. Apart from self-attention, we evaluate the entanglement between brain voxels to enhance connectivity exploration.
@@ 1 Introduction
Human perception of the world is determined by external stimuli and individual experiments that create brain activities involving non-linear interactions among 86 billion neuronal cells, which are thus complex [42, 52]. An effective method called functional Magnetic Resonance Imaging (fMRI) is usually applied to measure brain activities via Blood Oxygenation Level Dependent (BOLD) signals and implicitly has correlations or connectivities between neurons in response to stimuli [70, 79]. With the development of deep learning models, it is essential to learn the information of the measured fMRI signals with their corresponding visual data [61, 4].
@@
Vision-brain understanding aims to learn information from brain signals obtained from image stimuli. It offers the exploration interest of cognition and perception that potentially contributes to brain-computer interface [13] and beyond. Brain-to-image, especially fMRI-to-image, is one of the fundamental tasks in vision-brain understanding. Generative models including GANs [15] and diffusion models [53, 76] enable to reconstruct more realistic images [61, 44, 66, 58]. Moreover, cross-subject and multiple-subject brain decoding is taken into account [74, 59, 48]. Besides, image and brain retrievals are also essential to map the information of the brain signals and images into the same information space [27, 43, 58]. However, vision-brain understanding is still confronted with significant challenges that hinder its applications on a broader scale.
@@
Limitations in Prior Methods: Prior methods focus on the conditioned image generation models [66, 65] that ignore the information extraction in the brain signals. The other methods focus on the fMRI activity mapping to the embeddings of the images [58, 8, 48, 74]. However, these methods make assumptions about the roles of brain regions that hinder the exploration of complex information and connectivity in brain signals. Moreover, since the connectivities in the fMRI signals are crucial to extracting the information, the multi-layer perceptron (MLP) backbone models in prior methods [58, 59] show the missing correlation between fMRI voxels in the brain activities. Besides, self-attention is one of the conventional approaches to computing the correlation between elements [48, 8]. However, in fMRI voxels, the connectivities are based on the functional information of each brain region, while self-attention learns to find patterns for correlation computing which makes the model harder to learn and extract the connectivity. Fig. Quantum-Brain: Quantum-Inspired Neural Network Approach to Vision-Brain Understanding illustrates the difference between self-attention and the proposed approach.
@@
Motivations from Quantum Machine Learning: Far apart from traditional computers, quantum computers are emerging machines that perform quantum algorithms [2]. Quantum computing utilizes quantum theory to process data in quantum devices [20, 41] with existing models including quantum circuits, quantum annealing, and adiabatic quantum computation. It has been proved that quantum computers outperform classical computers in solving specific problems [16, 63]. In the noisy intermediate scale quantum (NISQ) era [47], while quantum computers cannot perform complex quantum algorithms for practical applications, the quantum computing theory provides a new mathematical formalism for computational tasks. Hence, quantum machine learning is introduced as a new learning paradigm utilizing quantum computation to enhance classical machine learning models [56, 5]. Moreover, quantum theory has been applied in classical algorithms and deep learning models, expected to improve computational performance and quality [14, 67].
@@
The quantum theory has two important properties, including superposition and entanglement. While the superposition can represent high-dimensional information, the entanglement can compute the connection between entities. Motivated by the quantum theory, we introduce a quantum-inspired network for vision-brain understanding. The proposed quantum-inspired network includes a new phase-shifting module to calibrate the value of the fMRI voxels, a voxel connection module to compute the connectivity between voxels via the entanglement, and a measurement-like projection module to project the information represented in the Hilbert space into the feature space. It is a simple but efficient method to compute the connectivity of the fMRI voxels for vision-brain understanding.
@@
Contributions of this Work: To contribute to the development of the vision-brain understanding, we introduce a novel Quantum-Brain approach, a quantum-inspired neural network, to learn the connectivity information in the brain signal corresponding to human perceptions. To the best of our knowledge, the proposed Quantum-Brain model is the first quantum-inspired approach to vision-brain understanding. The contributions of this work can be summarized as follows. First, motivated by the entanglement property in quantum computing, we proposed a novel Quantum-Inspired Voxel-Controlling Module for learning the connectivity in the brain signals. Second, a new Phase Shifting Module is introduced to calibrate the voxel values in the brain signals and enhance the robustness of the connectivity extraction. Third, as the current connectivity information is represented in the quantum space, a new Measurement-like Projection Module is presented to transform the extracted information into the feature space. Finally, through our experiments on the Natural Scene Dataset (NSD) benchmarks [1], including image-brain retrieval and fMRI-to-image reconstruction, our proposed method achieves state-of-the-art performance with the Top-1 accuracy of $95.1\%$ and $95.6\%$ on the image and brain retrieval benchmarks and Inception score of $95.3\%$ on the fMRI-to-image reconstruction benchmark. Our experimental results show the effectiveness of brain feature representation for vision-brain understanding.
@@ 2 Related Work
This section describes prior visual brain signal decoding methods and quantum-inspired methods for deep learning problems.
@@ 2.1 Visual Brain Signal Decoding
Researchers showed that visual information, such as spatial position [68], orientation [19, 23], and coarse image category [11, 18], could be decoded from fMRI signals using linear models. With the advent of generative adversarial networks [15], more advanced decoding methods became possible and enabled researchers to map brain activity to the latent space of these models to reconstruct images such as handwritten digits [55], human faces [72, 12], and natural scenes [61, 44, 60]. Recently, with the development of multimodal contrastive models like CLIP [49], diffusion models [21, 64] like Stable Diffusion [53], and new large-scale fMRI datasets [1], fMRI-to-image reconstructions have achieved an unprecedented level of quality [66, 43, 17].
Lin et al. [27] reconstructed seen images by aligning fMRI voxel data into CLIP space and generating outputs via a fine-tuned Lafite GAN [80]. Moreover, the alignment between the fMRI voxels and the CLIP space enabled the model to perform image-brain retrieval tasks. Ozcelik and VanRullen [43] employed a two-stage approach with Versatile Diffusion [76], utilizing low- and high-level processing. Gu et al. [17] also applied a similar pipeline and extended on Ozcelik et al. [44] by implementing IC-GAN [7] for reconstruction. Unlike prior methods, they did not flatten voxels but aligned them to SwAV [6] features using surface-based convolutional networks. Meanwhile, Takagi and Nishimoto [66] used ridge regression to align the fMRI voxels into Stable Diffusion latent and CLIP text patients by selecting different voxel regions for various components. MinD-Vis [8] addressed fMRI-to-image reconstruction by pre-training a masked brain encoder on a distinct large-scale fMRI dataset rather than the Natural Scenes Dataset [1], producing a more informative latent input for their image reconstruction model. Building on this, MinD-Video [9] expanded the MinD-Vis approach to reconstruct video instead of single images. MindEye [58] enhanced fMRI-to-image reconstruction and image-brain retrieval by separating the feature representation via contrastive learning and diffusion priors. Additionally, Psychometry [48], MindBridge [74], and MindEye2 [59] focused on a unified fMRI feature extraction from multiple subjects. Meanwhile, previous work analyzed the correlation between fMRI signals and deep learning models [77, 36, 37].
@@ 2.2 Quantum-Inspired Machine Learning
Recently, quantum computing has been applied to machine learning tasks, i.e., clustering [29, 35, 38, 40], principal component analysis [30], least-squares fitting [57, 24], and binary classification [51]. With the rise of quantum neural networks, variational quantum algorithms [45, 33] were introduced to alternate the classical deep learning algorithms [10, 3, 22, 54, 34, 39].
Meanwhile, quantum-inspired computing refers to practical methods derived from quantum computing concepts. The mathematical formalism of quantum computing has been adopted in various machine learning tasks, including natural language processing [25], computer vision [62, 67, 78], information retrieval [69, 71] and multimodal analysis [14, 26]. Van et al. [71] applied the quantum probability of object representation in information retrieval tasks. Li et al. [25] proposed a quantum-inspired network in language models for semantic matching to achieve better outcome quality and interpretability. Li et al. [26] investigated multimodal feature fusion methods via a quantum-inspired neural network for conversation emotion recognition. Meanwhile, Gkoumas et al. [14] introduced a video semantic recognition network by fusing multimodal information at the decision level. Moreover, quantum-inspired methods have also been developed for image-processing tasks. Shi et al. [62] presented complex-valued convolutional neural networks to process high-dimensional data and represent better non-linear space. Especially, Tang et al. [67] exploited the wave-like function to represent an image via phase-shifting operations to learn the relationship between patches. Meanwhile, Zhang et al. [78] utilize the quantum state representation to enhance the hyperspectral image feature extraction.
@@ 3 Preliminaries
Most quantum computers are constructed based on the quantum computing model [2, 41]. A quantum computation includes state initialization, unitary evolution, and measurement. The quantum device is first set to the initial state. The quantum state is modified by applying unitary operators to quantum devices. Then, the initial state of the system is transformed into the final state containing processing results in the quantum device. To obtain classical information from the quantum device, measurements are utilized to extract processing results from the final state.
[…]
@@
where $A_{n}$ is a probability amplitude satisfying $\sum_{n=1}^{N}|A_{n}|^{2}=1$, $i$ is the imaginary unit, and $\theta_{n}$ indicates the phase. A simple example of basis state vectors for $\mathcal{H}^{N}$ is the standard unit basis. In particular, the standard basis of $\mathcal{H}^{2}$ is formed by column vectors $|0\rangle_{2}=(1,0)^{\top}$ and $|1\rangle_{2}=(0,1)^{\top}$. Here, we use the notation as in Eqn. (2) to represent the standard basis of $\mathcal{H}^{N}$.
$|n\rangle_{N},n=0,1,\dots,N-1$ (2)
Respectively, $\langle n|_{N}$ denotes the conjugate transpose of $|n\rangle_{N}$. All components of $|n\rangle_{N}$ are 00 except the $n$-th component is $1$. In quantum theory, Eq. (1) represents the superposition of basis states.
A unitary operator $U$ denotes a quantum state evolution that transforms the initial states $|\psi\rangle$ to final states $|\psi^{\prime}\rangle$. These operators manipulate the phases $\theta_{n}$ and amplitudes $A_{n}$ by applying specific physical operations to the quantum system. In quantum computing, projection-valued measure (PVM) is a function that can collapse the system state from the superposition of multiple basis states to one basis state. The mathematical formulation of PVM is a set of projection operators $\{\Pi_{m}=|\phi_{m}\rangle\langle\phi_{m}|\}_{m=1}^{M}$. As described in the Born rule [41], $\Pi_{m}$ projects the quantum state $|\psi^{\prime}\rangle$ to the basis state $|\phi_{m}\rangle$ with a probability $P(|\phi_{m}\rangle)$ as in Eqn. (3).

### v2

@@ Abstract
Vision-brain understanding aims to extract semantic information about brain signals from human perceptions. Existing deep learning methods for vision-brain understanding are typically introduced within a traditional learning paradigm, lacking the ability to learn the connectivities between brain regions. Meanwhile, the quantum computing theory offers a new paradigm for designing deep learning models in classical systems. Motivated by the connectivities in brain signals and the entanglement properties in quantum computing, we propose a novel Quantum-Brain approach, a quantum-inspired neural network, to tackle the vision-brain understanding problem. To compute the connectivity between areas in brain signals, we introduce a new Quantum-Inspired Voxel-Controlling module to learn the impact of a brain voxel on others represented in the Hilbert space. To effectively learn connectivity, a novel Phase-Shifting module is presented to calibrate the value of the brain signals. Finally, we introduce a new Measurement-like Projection module to present the connectivity information from the Hilbert space into the feature space. The proposed approach can learn to find the connectivities between fMRI voxels and enhance the semantic information obtained from human perceptions. Our experimental results on the Natural Scene Dataset benchmarks demonstrate the effectiveness of the proposed method, achieving Top-1 accuracies of $95.5\%$ and $95.3\%$ for image and brain retrieval tasks, respectively, and an Inception score of $95.6\%$ for the fMRI-to-image reconstruction task. Our proposed quantum-inspired network introduces a novel paradigm for solving vision-brain problems using quantum computing theory.
Index Terms:
Quantum-inspired machine learning, quantum computing theory, classical neural networks, and human vision-brain.
Figure 1: Quantum-Brain for Vision-Brain Understanding. Motivated by the connectivities in brain signals and entanglement in quantum theory, we propose Quantum-Brain, a quantum-inspired neural network to explore the connectivities in brain signals for vision-brain understanding. Apart from self-attention, we evaluate the entanglement between brain voxels to enhance the exploration of connectivity.
@@ 1 Introduction
Human perception of the world is determined by external stimuli and individual experiments that create brain activities involving non-linear interactions among 86 billion neuronal cells, which are thus complex [1, 2]. An effective method called functional Magnetic Resonance Imaging (fMRI) is typically used to measure brain activity via Blood Oxygenation Level Dependent (BOLD) signals, implicitly revealing correlations or connectivities between neurons in response to stimuli [3, 4]. With the development of deep learning models, it is essential to learn the information contained in the measured fMRI signals, along with their corresponding visual data [5, 6].
@@
Vision-brain understanding aims to learn information from brain signals obtained from image stimuli. It offers the exploration interest of cognition and perception that potentially contributes to brain-computer interface [7] and beyond. Brain-to-image, especially fMRI-to-image, is one of the fundamental tasks in understanding the brain and vision. Generative models including GANs [8] and diffusion models [9, 10] enable to reconstruct more realistic images [5, 11, 12, 13]. Moreover, cross-subject and multiple-subject brain decoding is taken into account [14, 15, 16]. Besides, image and brain retrievals are also essential to map the information of the brain signals and images into the same information space [17, 18, 13]. However, understanding the vision-brain connection is still confronted with significant challenges that hinder its broader-scale applications.
@@
Limitations in Prior Methods: Prior methods focus on conditioned image generation models [12, 19] that overlook information extraction from brain signals. The other methods focus on the fMRI activity mapping to the embeddings of the images [13, 20, 16, 14]. However, these methods make assumptions about the roles of brain regions that hinder the exploration of complex information and connectivity in brain signals. Moreover, since the connectivities in the fMRI signals are crucial to extracting the information, the multi-layer perceptron (MLP) backbone models in prior methods [13, 15] show the missing correlation between fMRI voxels in the brain activities. Besides, self-attention is one of the conventional approaches to computing the correlation between elements [16, 20]. However, in fMRI voxels, the connectivities are based on the functional information of each brain region. At the same time, self-attention learns to find patterns for correlation computing, which makes the model harder to learn and extract the connectivity. Fig. 1 illustrates the difference between self-attention and the proposed approach.
@@
Motivations from Quantum Machine Learning: Far apart from traditional computers, quantum computers are emerging machines that perform quantum algorithms [21]. Quantum computing utilizes quantum theory to process data in quantum devices [22, 23] with existing models including quantum circuits, quantum annealing, and adiabatic quantum computation. It has been proved that quantum computers outperform classical computers in solving specific problems [24, 25]. In the noisy intermediate-scale quantum (NISQ) era [26]. In contrast, quantum computers cannot perform complex quantum algorithms for practical applications, the quantum computing theory provides a new mathematical formalism for computational tasks. Hence, quantum machine learning is introduced as a new learning paradigm utilizing quantum computation to enhance classical machine learning models [27, 28]. Moreover, quantum theory has been applied to classical algorithms and deep learning models, which are expected to improve computational performance and quality [29, 30]. Furthermore, prior clinical and theoretical physics studies [31, 32, 33] demonstrate the relationships between quantum theories and the human brain. Hence, our proposed work represents the first successful machine vision study to realize these hypotheses.
@@
The quantum theory has two important properties, including superposition and entanglement. While the superposition can represent high-dimensional information, the entanglement can compute the connection between entities. Motivated by quantum theory, we introduce a quantum-inspired network for understanding the vision-brain connection. The proposed quantum-inspired network includes a new phase-shifting module to calibrate the value of the fMRI voxels, a voxel connection module to compute the connectivity between voxels via entanglement, and a measurement-like projection module to project the information represented in the Hilbert space into the feature space. It is a simple yet efficient method for computing the connectivity of fMRI voxels in the context of vision-brain understanding.
@@
Contributions of this Work: To contribute to the development of a vision-brain understanding, we introduce a novel Quantum-Brain approach, a quantum-inspired neural network, to learn the connectivity information in brain signals corresponding to human perceptions. To the best of our knowledge, the proposed Quantum-Brain model is the first quantum-inspired approach to vision-brain understanding. The contributions of this work can be summarized as follows. First, motivated by the entanglement property in quantum computing, we proposed a novel Quantum-Inspired Voxel-Controlling Module for learning the connectivity in the brain signals. Second, a new Phase Shifting Module is introduced to calibrate the voxel values in the brain signals and enhance the robustness of the connectivity extraction. Third, as the current connectivity information is represented in the quantum space, a new Measurement-like Projection Module is presented to transform the extracted information into the feature space. Finally, through our experiments on the Natural Scene Dataset (NSD) benchmarks [34], including image-brain retrieval and fMRI-to-image reconstruction, our proposed method achieves state-of-the-art performance with the Top-1 accuracy of $95.5\%$ and $95.3\%$ on the image and brain retrieval benchmarks and Inception score of $95.6\%$ on the fMRI-to-image reconstruction benchmark. Our experimental results demonstrate the effectiveness of brain feature representation for understanding vision and the brain.
@@ 2.1 Visual Brain Signal Decoding
Researchers have shown that visual information, such as spatial position [35], orientation [36, 37], and coarse image category [38, 39], can be decoded from fMRI signals using linear models. With the advent of generative adversarial networks [8], more advanced decoding methods became possible and enabled researchers to map brain activity to the latent space of these models to reconstruct images such as handwritten digits [40], human faces [41, 42], and natural scenes [5, 11, 43]. Recently, with the development of multimodal contrastive models like CLIP [44], diffusion models [45, 46] like Stable Diffusion [9], and new large-scale fMRI datasets [34], fMRI-to-image reconstructions have achieved an unprecedented level of quality [12, 18, 47].
Lin et al. [17] reconstructed seen images by aligning fMRI voxel data into CLIP space and generating outputs via a fine-tuned Lafite GAN [48]. Moreover, the alignment between the fMRI voxels and the CLIP space enabled the model to perform image-brain retrieval tasks. Ozcelik and VanRullen [18] employed a two-stage approach, utilizing Versatile Diffusion [10] for both low- and high-level processing. Gu et al. [47] also applied a similar pipeline and extended on Ozcelik et al. [11] by implementing IC-GAN [49] for reconstruction. Unlike prior methods, they did not flatten voxels but aligned them to SwAV [50] features using surface-based convolutional networks. Meanwhile, Takagi and Nishimoto [12] used ridge regression to align the fMRI voxels into Stable Diffusion latent and CLIP text patients by selecting different voxel regions for various components. MinD-Vis [20] addressed fMRI-to-image reconstruction by pre-training a masked brain encoder on a distinct large-scale fMRI dataset rather than the Natural Scenes Dataset [34], producing a more informative latent input for their image reconstruction model. Building on this, MinD-Video [51] expanded the MinD-Vis approach to reconstruct video instead of single images. MindEye [13] enhanced fMRI-to-image reconstruction and image-brain retrieval by separating the feature representation via contrastive learning and diffusion priors. Additionally, Psychometry [16], MindBridge [14], and MindEye2 [15] focused on a unified fMRI feature extraction from multiple subjects. Meanwhile, previous work analyzed the correlation between fMRI signals and deep learning models [52, 53, 54]. Additionally, with the recent advancements in continual learning methods [55, 56] and domain generalization [57] across multiple modalities, COBRA [58] has studied vision-brain understanding in continual learning settings. Furthermore, a preliminary study of the human brain in quantum machine learning was introduced [59], showing the potential of the relationship between the human brain and quantum theory.
@@ 2.2 Quantum-Inspired Machine Learning
Recently, quantum computing has been applied to machine learning tasks, i.e., clustering [60, 61, 62, 63], principal component analysis [64], least-squares fitting [65, 66], binary classification [67], and quadratic optimization [68, 69, 70]. With the rise of quantum neural networks, variational quantum algorithms [71, 72] were introduced to replace the classical deep learning algorithms, including convolutional neural networks [73], recurrent neural networks [74], generative models [75, 76, 77, 78], reinforcement learning [79, 80], and mixture of experts [81, 82].
Meanwhile, quantum-inspired computing refers to practical methods derived from concepts in quantum computing. The mathematical formalism of quantum computing has been adopted in various machine learning tasks in classical systems, including natural language processing [83], computer vision [84, 30, 85], information retrieval [86, 87], and multimodal analysis [29, 88]. Van et al. [87] applied the quantum probability of object representation in information retrieval tasks. Li et al. [83] proposed a quantum-inspired network for language models, aiming to achieve better outcome quality and interpretability through semantic matching. Li et al. [88] investigated multimodal feature fusion methods via a quantum-inspired neural network for conversation emotion recognition. Meanwhile, Gkoumas et al. [29] introduced a video semantic recognition network by fusing multimodal information at the decision level. Moreover, quantum-inspired methods have also been developed for image-processing tasks. Shi et al. [84] presented complex-valued convolutional neural networks to process high-dimensional data and represent a better nonlinear space. Notably, Tang et al. [30] utilized a wave-like function to represent an image through phase-shifting operations, thereby learning the relationship between patches. Meanwhile, Zhang et al. [85] utilize the quantum state representation to enhance the extraction of hyperspectral image features.
@@ 3 Preliminaries
Most quantum computers are constructed based on the quantum computing model [21, 23]. A quantum computation includes state initialization, unitary evolution, and measurement. The quantum device is first set to the initial state. The quantum state is modified by applying unitary operators to quantum devices. Then, the initial state of the system is transformed into the final state, which contains the processing results in the quantum device. To obtain classical information from the quantum device, measurements are utilized to extract processing results from the final state.
Mathematically, a quantum state can be represented by a state vector in an $N$-dimensional Hilbert space $\mathcal{H}^{N}$. We adopt the Dirac notations in quantum theory to denote the quantum states [23]. The notations $|\cdot\rangle$ and $\langle\cdot|$ denote a Ket to represent a complex-valued unit vector and a Bra to indicate its conjugate transpose. In detail, given a set of basis state vectors $\{|\phi_{n}\rangle\}^{N}_{n=1}$, a quantum state $|\psi\rangle$ can be represented as a combination of basis state vectors as shown in Eqn. (1).
@@
where $A_{n}$ is a probability amplitude satisfying $\sum_{n=1}^{N}|A_{n}|^{2}=1$, $i$ is the imaginary unit, and $\theta_{n}$ indicates the phase. A simple example of basis state vectors for $\mathcal{H}^{N}$ is the standard unit basis. In particular, the standard basis of $\mathcal{H}^{2}$ is formed by column vectors $|0\rangle_{2}=(1,0)^{\top}$ and $|1\rangle_{2}=(0,1)^{\top}$. Here, we use the notation $|n\rangle_{N}$ to represent the standard basis of $\mathcal{H}^{N}$. Respectively, $\langle n|_{N}$ denotes the conjugate transpose of $|n\rangle_{N}$. All components of $|n\rangle_{N}$ are 0 except the $n$-th component is $1$. In quantum theory, Eqn. (1) represents the superposition of basis states.
A unitary operator $U$ denotes a quantum state evolution that transforms the initial states $|\psi\rangle$ to final states $|\psi^{\prime}\rangle$. These operators manipulate the phases $\theta_{n}$ and amplitudes $A_{n}$ by applying specific physical operations to the quantum system. In quantum computing, a projection-valued measure (PVM) is a function that can collapse the system state from a superposition of multiple basis states to a single basis state. The mathematical formulation of PVM is a set of projection operators $\{\Pi_{m}=|\phi_{m}\rangle\langle\phi_{m}|\}_{m=1}^{M}$. As described in the Born rule [23], $\Pi_{m}$ projects the quantum state $|\psi^{\prime}\rangle$ to the basis state $|\phi_{m}\rangle$ with a probability $P(|\phi_{m}\rangle)$ as in Eqn. (2).